OpenStudy (toxicsugar22):

You add 9.3g of iron to 26.60mL of water and observe that the volume of iron and water together is 27.78mL . Calculate the density of iron.

OpenStudy (ankit042):

volume = mass/density density of water is 1 g/mL 9.3/d +26/1 = 27.78

OpenStudy (toxicsugar22):

241.8

OpenStudy (toxicsugar22):

i multiplied 9.3 by 26 is that right

OpenStudy (ankit042):

9.3/d = (27.78-26)

OpenStudy (toxicsugar22):

1.78

OpenStudy (toxicsugar22):

16.554

OpenStudy (toxicsugar22):

16.554 is the anser right

OpenStudy (ankit042):

NO 9.3/1.78

OpenStudy (toxicsugar22):

5.2224

OpenStudy (toxicsugar22):

right

OpenStudy (toxicsugar22):

can you still hep me